OpenStudy (anonymous):

What is the solution to 5/(x^2 + 3x + 2) - 1/(x + 2) = 1/(3x + 3) ?

OpenStudy (anonymous):

5/2 do you want me to explain

OpenStudy (anonymous):

id rather just have the answer. Last time I got all confused.

OpenStudy (anonymous):

medal please

OpenStudy (campbell_st):

rewrite the problem \[\frac{5}{(x +2)(x +1)} - \frac{1}{(x+2)} = \frac{1}{3(x +1)}\]

OpenStudy (campbell_st):

multiply every term by (x +2)(x + 1) then \[5 - (x +1) = \frac{x+2}{3}\] multiply both sides by 3 \[15 - 3(x +1) = (x + 2)\] now solve for x